OpenStudy (anonymous):

Find the equation of the tangent line to the graph f(x)=x^2-2at x=xnaught

OpenStudy (phi):

can you find the derivative of f(x) with respect to x ? the derivative will be a function that defines the slope of the tangent line at any given x Evaluate f'(x) at \( x_0 \) to get m, the slope of the tangent line. Evaluate f(x) at \( x_0 \) to get \( y_0 \) Once you have slope m and a point on the line \( ( x_0, y_0) \), you can find the equation of the line using \[ y - y_0 = m(x-x_0) \]

OpenStudy (anonymous):

y=x^2 - 2 dy/dx = 2x You know the equation of the line when a point and slope are given: y-yo = m(x-xo) Plugging in the value, we get y-xo^2 - 2 = 2xo(x-xo) y = 2xo*x - (xo^2+2) The above equation is in y = mx+c form.